Operations Bulletin 020315

From GridPP Wiki
Jump to: navigation, search

Bulletin archive

Week commencing 24th February 2015
Task Areas
General updates

Tuesday 24th February

Tuesday 17th February

  • There was a GDB last week. See the agenda and minutes.
    • Introduction
    • Sys-on-Chip
    • Ops coordination
    • pre-GDB on cloud traceability
    • HSF workshop
    • INDIGO datacloud
  • Those needing an LCG associated CERN account should reply to Jeremy's mail to TB-SUPPORT (12/02/14).
  • The final EGI-InSPIRE review took place last week. Feedback so far is very positive. EGI.eu is now working on its 2015-2020 strategy and 'open commons' is playing a large role.
  • Adam Huffman has left Imperial. Thanks to Adam for his contributions.
  • Working with DiRAC ...
  • There was a GridPP technical meeting last Friday. Any minutes?
  • GridPP: Low availability/reliability sites may be contacted regarding options for adopting a lightweight setup. Sites not meeting experiment needs will be reviewed against suitability for Tier-2 inclusion - note ATLAS focus is moving to sites in the 80%-90% ASAP region.
WLCG Operations Coordination - Agendas

Monday 23rd February

  • There was an ops coordination meeting last Thursday 19th. Agenda. Minutes.
  • News: Following up on survey responses - some misunderstandings. Savannah closed.
  • Baselines: UMD 3.11.0 released (APEL-client/SSM, CREAM, Gfal-utilis and DPM 1.8.9 updates.
  • MW Issues: ARGUS failures fixed in PAP 1.6.4. LHCb DPM Brazil site checksum issue. FTS3 shares not respected.
  • T0&1 services: CERN EOS Aquamarine for ALICE and LHCb. Various dacha upgrades.
  • T0 news: LHCb LFC migrating to DFC. Then LFC now only used by ops. VOMRS final date 2nd March.
  • T1: NTR
  • T2: NTR
  • ALICE: Normal to high activity. gilte VOBox SSL errors. Moved to WLCG VOBoxes.
  • ATLAS: High loads. Some MC improvement possible - job lengths 100->200 events. MC15a at final validation (will last 3 months). Taking cosmics. FTS3 shows new bugs (intra-VO shows). SARA data reproduction ongoing.
  • CMS: Taking cosmics. Some issues with T0 VMs. Run2 MC campaign at T2s. Considering to drop CRC32 checksum (just use Adler). Tape staging tests ongoing.
  • LHCb: Finalising Run1 legacy stripping. Migrating LFC to DFC (w/c 23).
  • HTTP deployment task force: Mandate to ID functionality set expts. require; provide recipes & recommendations to sites; coordinate deployment. Action to find out experiment desire for http.
  • glexec (in Panda): Debugging RAL & RALPP. Big sites in. 54 covered.
  • SHA-2: VOMS configuration reminders will go out when VOMRS stopped. (2/03)
  • MJF: NTR
  • MW Readiness: Thanks to those contributing. More needed to install Package Reporter. Next meeting 18th March 16:00 CET.
  • MultiCore: NTR
  • IPv6: NTR
  • Squid mon & HTTP proxies: NTR
  • Network & Transfer metrics: NTR.

Tuesday 17th February

Tier-1 - Status Page

Tuesday 24th February

  • There were two brief network outages at RAL this morning (24th) while the core network team rebooted a core switch.
  • Some racks in the datacentre are being moved today to make way for latest procurement. This will affect one generation of production diskservers. After consultation with the VOs, we have declared an outage in the GOCDB for Atlas, LHCb and GEN. CMSTape is affected but CMS have elected to stay up.
  • The problems with our primary network router are still being followed up. The vendor has been given data from the last failover test.
Storage & Data Management - Agendas/Minutes

Monday 23rd February

  • From February 17th expect xrootd version 4 will be added to the EPEL repository. Update not transparent if xrootd supported.

Tuesday 17th February

  • Chris suggests linking the WebDav page from the storage wiki.

Wedn 11 Feb

Tuesday 10th February

Accounting - UK Grid Metrics HEPSPEC06 Atlas Dashboard HS06

Tuesday 24th February

  • UKI-SOUTHGRID-SUSX now highlighted as not publishing for 3 months.
Documentation - KeyDocs

See the worst KeyDocs list for documents needing review now and the names of the responsible people.

Tuesday 24th February

Tuesday 17th February

  • KeyDocs review continues. Deadline for removals requests 25th February.

Tuesday 10th February

  • Survey of KeyDocs has started. If you own a document please take note. Documents remaining after this exercise will need to be brought fully up-to-date.
  • A reminder that Steve created this Current Activities capture page. Only ARGUS at the moment!?
  • See also the documents from Tom linked in the 'Other VOs' section.

Tuesday 3 Feb

  • New section in Wiki called "Project Management Pages".
The idea is to cluster all Self-Edited Site Tracking Tables
in here. Sites should keep entries in Current Activities
up to date. Once a Self-Edited Site Tracking Tables has
served its purpose, PM to move it to  Historical Archive 
or otherwise dispose of the table.
Interoperation - EGI ops agendas

Monday 9th February

  • The agenda for February's EGI ops meeting is here.
    • Argus PAP update (1.6.2) out imminently. This should fix the SSL problems for sites affected after a java update. Hopefully this will solve Raul's problems in his ticket 111505.
    • Other update notifications include APEL 1.4.0 (with a bunch of new features), Storm 1.11.6, UNICORE v7.2.0 and a new version of Globus in EPEL testing. UMD release 3.11.0 being prepared.
    • Multicore republishing was discussed - if sites need to republish multicore data previously parsed as single core then they need to reparse the logs. Any site wanting to republish is asked to submit a ticket to apel support first.
    • Operational issues were discussed whereby sites saw problems running ARGUS and CREAM on the same node, special care is taken if you want to take this mad path. See tickets 111464 and 111373.
    • The decommissioning of the EMI2 APEL service and the UMD repositories was mentioned.
    • Sites are advised to avoid using MySQL v.5.0 where possible.
    • A comment was passed that the EPEL5 repository was becoming harder to keep up to date, bringing the effective EOL of SL5 closer for the community
      • SL7 is a topic that still needs to be discussed in earnest in this forum.
    • Next meeting is March 9th - effort to keep the number of meetings down.

Monitoring - Links MyWLCG

Monday 7th December

On-duty - Dashboard ROD rota

Monday 23rd February

  • Another quiet week.
  • Gordon -> Andrew

Monday 16th February

  • Uneventful week. Nothing to report to ops meeting.
Rollout Status WLCG Baseline

Tuesday 17th February

  • A message from Maria Dimou: A request to you and the sites you collaborate with, to please install the MW Package Reporter.

The current version, documented in the wiki , it is very easy to install and addresses all site and security requirements.

  • Quick check on: 1. ARC CE readiness testing. 2. Machine job features testing.

Tuesday 20th January

  • From Cristina's GDB talk last week note that EMI repositories will be frozen and the product team releases will become UMD-preview (repository content and webpages similar but now managed).


Security - Incident Procedure Policies Rota

Tuesday 24th February

  • Tracking approvals in GOCDB.
  • EGI SVG Advisory - dCache vulnerability for some access methods [SVG EGI-SVG-2015-8183]
  • New security officer starts this week!

Tuesday 17th February

  • The IGTF is about to release (23/2) an update to the trust anchor repository (1.62)

Services - PerfSonar dashboard | GridPP VOMS

- This includes notifying of (inter)national services that will have an outage in the coming weeks or will be impacted by work elsewhere. (Cross-check the Tier-1 update).

Tuesday 3rd February

Tuesday 20th January


Monday 23rd February 2015, 15.00 GMT
Only 15 Open UK tickets this week. Feel free to bring up any ticket-based issues of your own on this quiet week.

This CMS glexec hammercloud ticket is looking a little quiet - no update for a while. If it's continuing offline or waiting on input could it at least be put On Hold? In progress (11/2) (The Tier-1 version of this ticket, 111699, seems to be chugging along fine - there might be useful snippets in there).

The Cloud accounting probe ticket was reopened, asking if 100IT ticketed apel support (I assume contacting them via other means would work too) otherwise the new cloud accounting won't be properly republished. Reopened (20/2)

IMPERIAL (but not really their issue)
Tom opened a ticket after another cern@school user had troubles using the IC SE - there has been some problems with the newer versions of the dirac UI. Sometimes it's better to go Vintage! Although after trying Simon and Daniela haven't been able to reproduce the failure - perhaps something's up with the user's UI? Waiting for reply (23/2)

Sussex's Perfsonar ticket. I know Matt RB has put the ticket On Hold and is very busy - is there any news/anything we can do to help? On Hold (21/1

Sno+ gfal copy problems at RAL. Brian informs us that the latest version of the gfal tools works for him and has asked if they work for Matt M. and Co. Did you get these packages out of epel/epel-testing or somewhere else Brian? Waiting for reply (18/2)

Tools - MyEGI Nagios

Tuesday 17th February

  • Another period where message brokers were temporarily unavailable seen yesterday. Any news on the last follow-up?

Tuesday 27th January

  • Unscheduled outage of the EGI message broker (GRNET) caused a short-lived disruption to GridPP site monitoring (jobs failed) last Thursday 22nd January. Suspect BDII caching meant no immediate failover to stomp://mq.cro-ngi.hr:6163/ from stomp://mq.afroditi.hellasgrid.gr:6163/

VOs - GridPP VOMS VO IDs Approved VO table

Tuesday 17th February Two changes to approved VOs (https://www.gridpp.ac.uk/wiki/GridPP_approved_VOs)

  • LSST uses port 15003 (had been 15002, clashing with dzero)
  • t2k has included a note that it';s software is now distrinuted via CVMFS.

Tuesday 17th February

  • Some interest from UK LIGO users. Catalin setting up CVMFS. Issue with VOMS. "Our users have access to CILogon basic CA certificates, which I understand can in principal be entered into VOMS to provide a VO identity."

Tuesday 10th February

  • There is now a Quick Guide to CVMFS.
  • Part of Tom's quick guide series! There is also a Quick Guide to DIRAC.
  • Feedback on both welcome... on the CVMFS front some mapping issues for regional VOs are being examined.
  • GalDyn (UCLan galaxy simulations): test jobs have been submitted via DIRAC; now moving onto data storage testing via the official DIRAC tutorials.
  • CERN@school has its first student on the grid with a grid certificate! LUCID data arriving.
  • The CERN VM mechanism looks like it could provide an interesting way for new users to get a grid-ready UI by installing VirtualBox (or similar), downloading the CERN VM image, and applying a GridPP-specific context.

Site Updates

Tuesday 24th February

  • Next review of status today.

Tuesday 27th January

  • Squids not in GOCDB for: UCL; ECDF; Birmingham; Durham; RHUL; IC; Sussex; Lancaster
  • Squids in GOCDB for: EFDA-JET; Manchester; Liverpool; Cambridge; Sheffield; Bristol; Brunel; QMUL; T1; Oxford; Glasgow; RALPPD.

Tuesday 2nd December

  • Multicore status. Queues available (63%)
    • YES: RAL T1; Brunel; Imperial; QMUL; Lancaster; Liverpool; Manchester; Glasgow; Cambridge; Oxford; RALPP; Sussex (12)
    • NO: RHUL (testing); UCL; Sheffield (testing); Durham; ECDF (testing); Birmingham; Bristol (7)
  • According to our table for cloud/VMs (26%)
    • YES: RAL T1; Brunel; Imperial; Manchester; Oxford (5)
    • NO: QMUL; RHUL; UCL; Lancaster; Liverpool; Sheffield; Durham; ECDF; Glasgow; Birmingham; Bristol; Cambridge; RALPP; Sussex (14)
  • GridPP DIRAC jobs successful (58%)
    • YES: Bristol; Glasgow; Lancaster; Liverpool; Manchester; Oxford; Sheffield; Brunel; IC; QMUL; RHUL (11)
    • NO: Cambridge; Durham; RALPP; RAL T1 (4) + ECDF; Sussex; UCL; Birmingham (4)
  • IPv6 status
    • Allocation - 42%
    • YES: RAL T1; Brunel; IC; QMUL; Manchester; Sheffield; Cambridge; Oxford (8)
    • NO: RHUL; UCL; Lancaster; Liverpool; Durham; ECDF; Glasgow; Birmingham; Bristol; RALPP; Sussex
  • Dual stack nodes - 21%
    • YES: Brunel; IC; QMUL; Oxford (4)
    • NO: RHUL; UCL; Lancaster; Glasgow; Liverpool; Manchester; Sheffield; Durham; ECDF; Birmingham; Bristol; Cambridge; RALPP; Sussex, RAL T1 (15)

Tuesday 21st October

  • High loads seen in xroot by several sites: Liverpool and RALT1... and also Bristol (see Luke's TB-S email on 16/10 for questions about changes to help).

Tuesday 9th September

  • Intel announced the new generation of Xeon based on Haswell.

Meeting Summaries
Project Management Board - MembersMinutes Quarterly Reports


GridPP ops meeting - Agendas Actions Core Tasks


RAL Tier-1 Experiment Liaison Meeting (Wednesday 13:30) Agenda Meeting takes place on Vidyo.

Wednesday 25th February 2015

  • Operations report
  • Test of problematic Tier1 network router made last (Thursday) morning. We remain in contact with the vendor.
  • A Castor namesever box has been set-up to enable queries against Castor metadata to be made without affecting the throughput of production work.
  • A system has been set-up to provide Atlas with Castor information that is not supplied by the SRM.
WLCG Grid Deployment Board - Agendas MB agendas


NGI UK - Homepage CA


UK ATLAS - Shifter view News & Links

Atlas S&C week 2-6 Feb 2015


• Prodsys-2 in production since Dec 1st

• Deployment has not been transparent , many issued has been solved, the grid is filled again

• MC15 is expected to start soon, waiting for physics validations, evgen testing is underway and close to finalised.. Simulation expected to be broadly similar to MC14, no blockers expected.


• Rucio in production since Dec 1st and is ready for LHC RUN-2. Some fields need improvements, including transfer and deletion agents, documentation and monitoring.

Rucio dumps available.

Dark data cleaning

files declaration . Only Only DDM ops can issue lost files declaration for now, cloud support needs to fill a ticket.

• Webdav panda functional tests with Hammercloud are ongoing


Main page

DDM Accounting




• ASAP (ATLAS Site Availability Performance) in place. Every 3 months the T2s sites performing BELOW 80% are reported to the International Computing Board.





  • N/A
To note

  • N/A